OpenStudy (anonymous):

find the angle between <-5,-4> and <-4,-3>

OpenStudy (anonymous):

Let a ⃗=-5i-4j b ⃗=-4i-3j |a ⃗ |=√((-5)^2+(-4)^2 )=√41 |b ⃗ |=√((-4)^2+(-3)^2 )=5 Let 𝜃 be the angle between the vectors , then a ⃗.b ⃗=ab cos 𝜃=5 √41 cos 𝜃 Also a ⃗.b ⃗ =(-5i-4j).(-4i-3j)=20+12=32 Hence 32=5√41 cos 𝜃 Cos 𝜃=32/(5√41) 𝜃=cos^(-1)⁡〖32/(5√41)〗